OpenStudy (anonymous):

What is the sum of the arithmetic sequence 3, 9, 15…, if there are 22 terms?

OpenStudy (lgbasallote):

find A_22 first \[A_{22} = 3 + (22 - 1)6\] \[A_{22} = 3 + 126 = 129\] now that you have the last term solve for sum \[S_n = \frac{n(a_n + a_1)}{2}\] \[S_n = \frac{22(129 + 3)}{2}\] can you do it from there?

OpenStudy (lgbasallote):

btw...the formula for A_n is \[A_n = A_1 + (n-1)d\]